[Brahms-l] Shift report 20050524 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 24 2005 - 00:04:13 EDT
Shifter leader was kh
Shifters were og


Machine studies until about 22:30.
We changed the magnet polarities from B to A and moved the forward spectrometer from 2.3 to 4 deg.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 24 00:04:23 2005

This archive was generated by hypermail 2.1.8 : Tue May 24 2005 - 00:04:31 EDT